Circular No. 8263

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 159P/2003 UD_{16} (LONEOS) Images of comet P/2003 UD_{16} (cf. IAUC 8248) were identified and

measured by M. Meyer from Palomar Sky Survey photographs taken on 1989 Dec. 17 and 1991 Feb. 19. The resulting orbital elements by B. G. Marsden, given below, are taken from MPEC 2004-A18.

Epoch = 1989 Nov. 10.0 TT

T = 1989 Oct. 21.97484 TT  ω = 3.^{o}89285
e = 0.3839794  Ω = 56.46319 }2000.0
q = 3.6475040 AU  i = 23.26117

a = 5.9210751 AU n^{o} = 0.06840742 P = 14.41 years Epoch = 2004 Mar. 16.0 TT

T = 2004 Mar. 3.07823 TT  ω = 4.^{o}99908
e = 0.3806606  Ω = 55.16130 }2000.0
q = 3.6508526 AU  i = 23.42153

a = 5.8947523 AU n^{o} = 0.06886614 P = 14.31 years

COMETS C/2002 W9, W10, W11, W12, W13, W14, W15, W16, W17

Further to IAUC 8179, K. Battams has reported measurements of Kreutz sungrazing comets found by D. Evans (C/2002 W9), X. Leprette (C/2002 W10 and W15), X.-m. Zhou (C/2002 W11 and W13), J. Sachs (C/2002 W12), and S. Hönig (C/2002 W14, W16, and W17) on SOHO website images taken with the C2 coronagraph. C/2002 W9 and W12 were also observed in the C3 instrument. The reduced astrometry and orbital elements by B. G. Marsden appear on MPEC 2004-A03 and 2004-A05.

Comet  2002 UT  α_{2000}  δ_{2000}
C/2002 W9 Nov. 22.696  15^{h}44.^{m}7  −23^{o}37′
C/2002 W10  24.396  15 58.9  −22 13
C/2002 W11  25.079  16 02.1  −22 21
C/2002 W12  27.346  16 01.0  −26 19
C/2002 W13  27.963  16 15.4  −22 55
C/2002 W14  28.713  16 18.6  −23 06
C/2002 W15  29.064  16 21.6  −23 15
C/2002 W16  29.204  16 23.1  −23 22
C/2002 W17  29.288  16 20.7  −23 12
2004 January 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green